Strontium hydroxide is used in refining beet sugar and separating cyrstallizable sugar from molasses. It is commonly fo...

  1. Strontium hydroxide is used in refining beet sugar and separating cyrstallizable sugar from molasses. It is commonly found as the octahydrate, Sr(OH)2.8H2O. If a solution is prepared by dissolving 0.45 grams of this compound in sufficient water to make 100 ml of solution, determine the [H+], the [OH-], the pH and the pOH of the solution.

  1. Ethylamine (C2H5NH2) is a flammable liquid with an ammonia-like odor that is miscible with water. It is irritating to skin, mucous membranes and to the respiratory tract. The density of ethylamine is 0.689 g/ml. If the Ka of this C2H5 NH2 is 2 x 10-11, determine the [H+], the [OH-], the pH and the pOH of a solution prepared by dissolving 25.0 ml of this compound in 500 ml of solution.

  1. Phenylpropionic acid, C9H10O2, also known as hydrocinnamic acid , is a white crystalline solid with a sweet floral scent at room temperature. It has a wide variety of uses including cosmetics, food additives, and pharmaceuticals. If a 0.075 M solution of this acid has a pH of 2.89, determine the Ka for the acid.
Answer #1

roak 100 DOO Vinml molarmow 1) molarity of Solo192.840 = E 0145 & 1000 265.63 TOO E O DIGAM COH ? = 2*0.0169 = 0.0338 M [44]3) pH = 2.89 Shpe] = 10 PH -2.89 10 0.001288 M CI HO O + 40 = Cquqe 2 + H2B At equilibrilem. ka= [cqHqg0] [40] Cahol = 0.07